# Dense Hopfield Networks in the Teacher-Student Setting

### v1

@@ Appendix C Teacher-student replicated partition function
$Z=\sum_{\sigma^{a}}\exp\left(\beta\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right)$
@@
From this expression, we define the partition function of the inverse model to be $\mathcal{Z}=\sum_{\xi}\exp\left(-\beta H\left[\xi|\sigma\right]\right)$ (again, see Appendix A for the definition of $H\left[\xi|\sigma\right]$). The $L^{\text{th}}$ power of $\mathcal{Z}$ and its average then take the form
$\mathcal{Z}^{L}$ $=\sum_{\xi}\prod_{b}\exp\left(\beta\frac{p!}{N^{p-1}}\sum_{a}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right)$
@@
$*\displaystyle\approx\frac{1}{2^{MN}}\frac{1}{\left\langle\mathcal{Z}^{*}\right\rangle}\sum_{\xi^{*}}\sum_{\sigma}\exp\Bigg{(}\beta^{*}N\sum_{a\in\Gamma_{*}}\left[\frac{1}{N}\sum_{i}\xi^{*}_{i}\sigma^{a}_{i}\right]^{p^{*}}$
@@ Appendix D Teacher-student free entropy
$\frac{\log\left(Z^{*}\right)}{N}$ =$=\begin{cases}\left(-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}\right)\frac{1}{N}+\log 2&\quad p^{*}=2\\ \frac{1}{2}\left[\beta^{*}\right]^{2}\frac{p^{*}!}{N^{p^{*}-1}}+\log 2+\mathcal{O}\left(\frac{1}{N^{3p^{*}/2-2)}}\right)&\quad p^{*}\geq 3,\end{cases}$
@@
=$=\begin{cases}\frac{1}{2}\left(-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}\right)\alpha^{*}+\log 2&\quad p^{*}=2\\ \frac{1}{2}\left[\beta^{*}\right]^{2}\alpha^{*}+\log 2+\mathcal{O}\left(\frac{1}{N^{p^{*}/2-1}}\right)&\quad p^{*}\geq 3,\end{cases}$
@@
$\log\Bigg{\langle}\exp\Bigg{\{}\beta^{*}\frac{p^{*}!}{N^{p^{*}-1}}\sum_{i_{1}<...<i_{p^{*}}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p^{*}}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p^{*}}}$
$\quad+\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\Bigg{\}}\Bigg{\rangle}$
$\approx\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\sum_{b\neq c}\sum_{i_{1}<...<i_{p}}\sum_{j_{1}<...<j_{p}}\xi^{b}_{i_{1}}\xi^{c}_{j_{1}}...\xi^{b}_{i_{p}}\xi^{c}_{j_{p}}\left\langle\sigma^{a}_{i_{1}}\sigma^{a}_{j_{1}}...\sigma^{a}_{i_{p}}\sigma^{a}_{j_{p}}\right\rangle$
$\quad+\beta^{*}\beta\frac{p^{*}!}{N^{p^{*}-1}}\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p^{*}}}\sum_{j_{1}<...<j_{p}}\left\langle\xi^{*}_{i_{1}}\sigma^{a}_{i_{1}}...\xi^{*}_{i_{p^{*}}}\sigma^{a}_{i_{p^{*}}}\xi^{b}_{j_{1}}\sigma^{a}_{j_{1}}...\xi^{b}_{j_{p}}\sigma^{a}_{j_{p}}\right\rangle$
$\quad+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN+\frac{1}{2}\left[\beta^{*}\right]^{2}\frac{p^{*}!}{N^{p^{*}-1}}N.$
@@
$\log\Bigg{\langle}\exp\Bigg{\{}\beta^{*}\frac{p^{*}!}{N^{p^{*}-1}}\sum_{i_{1}<...<i_{p^{*}}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p^{*}}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p^{*}}}$
$\quad+\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\Bigg{\}}\Bigg{\rangle}$
$=\beta^{2}\frac{p!}{N^{p-1}}N\sum_{b<c}\left[\frac{1}{N}\sum_{i}\xi^{b}_{i}\xi^{c}_{i}\right]^{p}+\beta^{*}\beta\frac{p!}{N^{p-1}}N\sum_{b}\left[\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi^{b}_{i}\right]^{p}$
$\quad+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN+\frac{1}{2}\left[\beta^{*}\right]^{2}\frac{p!}{N^{p-1}}N,$
@@
$\approx\log\Bigg{[}\left\langle\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
$\quad\left\langle\exp\left\{\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right\}\right\rangle$
$\quad+\left\langle\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{j_{1}<...<j_{p}}\xi^{b}_{j_{1}}...\xi^{b}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{j_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle\Bigg{]},$
@@
$\log\left\langle\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
$\approx-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}$
@@
$\approx\beta^{2}\frac{p!}{N^{p-1}}N\sum_{b<c}\left[\frac{1}{N}\sum_{i}\xi^{b}_{i}\xi^{c}_{i}\right]^{p}+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN$
@@
$=\left\langle\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{j_{1}<...<j_{p}}\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\xi^{b}_{j_{1}}...\xi^{b}_{j_{p}}\ \xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{j_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
$=\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{j_{1}<...<j_{p}}\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\xi^{b}_{j_{1}}...\xi^{b}_{j_{p}}\left\langle\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{j_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
@@
$=\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}$
$\quad\left\langle\frac{p!}{N^{p}}\sum_{j_{1}<...<j_{p}}\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{i_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
$=V\left(\beta^{*},p\right)\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}},$
@@
$\approx-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}+\beta^{2}\frac{p!}{N^{p-1}}N\sum_{b<c}\left[\frac{1}{N}\sum_{i}\xi^{b}_{i}\xi^{c}_{i}\right]^{p}+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN$
$\quad+\left[1-2\beta^{*}\right]^{1/2}\exp\left(\beta^{*}\right)V\left(\beta^{*},p\right)\beta N\sum_{b}\left[\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi^{b}_{i}\right]^{p}.$
@@
$=\frac{\left[\beta^{*}\right]^{p/2}}{(p/2)!}\frac{p!}{N^{p/2}}$
@@
$\lambda$ $=\frac{\left[\beta^{*}\right]^{p/2}}{(p/2)!}N^{p/2-1}\sim\mathcal{O}\left(1\right)$
@@
$=\frac{1}{N^{4}}\sum_{j_{1}\neq j_{2}\neq j_{3}\neq j_{4}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\xi^{*}_{j_{3}}\xi^{*}_{j_{4}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\sigma^{a}_{j_{3}}\sigma^{a}_{j_{4}}$
$=\frac{1}{N^{4}}\left[\sum_{j_{1}\neq j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]\left[\sum_{j_{3}\neq j_{4}}\xi^{*}_{j_{3}}\xi^{*}_{j_{4}}\sigma^{a}_{j_{3}}\sigma^{a}_{j_{4}}\right]-\frac{4}{N^{3}}\left[\sum_{j_{1}\neq j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]-\frac{2}{N^{2}}$
$=\frac{1}{N^{2}}\left[\frac{2}{N}\sum_{j_{1}<j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]^{2}-\frac{4}{N^{2}}\left[\frac{2}{N}\sum_{j_{1}<j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]-\frac{2}{N^{2}}$
@@
$\frac{\partial\log\left(z^{*}\right)}{\partial\beta^{*}}$ $=\frac{\partial}{\partial\beta^{*}}\left[-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}\right]=\frac{2\beta^{*}}{1-2\beta^{*}}$
@@
$=\prod_{b}\sum_{\xi^{b}_{i}}\exp\left\{\beta k\xi^{b}_{i}\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}\xi^{b}_{i}+x\beta\sqrt{\alpha r}\xi^{b}_{i}\right\}$
$=\prod_{b}\left[2\cosh\left(\beta k\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}+x\beta\sqrt{\alpha r}\right)\right]$
$=2^{L}\cosh^{L}\left(\beta k\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}+x\beta\sqrt{\alpha r}\right)$
@@
$\exp\Bigg{\{}\beta k\sum_{b}\xi^{b}_{i}\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\sum_{b}\xi^{*}_{i}\xi^{b}_{i}+x\beta\sqrt{\alpha r}\sum_{b}\xi^{b}_{i}\Bigg{\}}\Bigg{\rangle}\Bigg{]}\Bigg{)}\Bigg{\}}$
@@
$\quad\cosh^{L}\left(\beta k\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}+x\beta\sqrt{\alpha r}\right)\Bigg{\rangle}\Bigg{]}\Bigg{)}\Bigg{\}}.$
@@
$f$ $*\displaystyle=\operatorname*{Extr}_{m,k,q,r,q^{*},r^{*}}\Bigg{\{}\beta^{*}\beta\alpha\left[q^{*}\right]^{p}-\frac{1}{2}\beta^{2}\alpha q^{p}+\beta m^{p}-\beta^{*}\beta\alpha r^{*}q^{*}$
@@
$f$ $=\operatorname*{Extr}_{m,k,q^{*},r^{*}}\Bigg{\{}\beta\eta\alpha\left[q^{*}\right]^{p}-\beta m^{p}-\beta\eta\alpha r^{*}q^{*}-\beta mk+\log 2$
$\quad+\Big{\langle}\log\left[\cosh\left(\beta\left[\eta\alpha r^{*}+kz\right]\right)\right]\Big{\rangle}_{z}\Bigg{\}}.$

### v2

@@ Appendix C Teacher-student replicated partition function
$Z=\sum_{\sigma^{a}}\exp\left(\beta\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right)\,,$
@@
Therefore, we define the partition function of the inverse model to be $\mathcal{Z}=\sum_{\xi}\exp\left(-\beta H\left[\xi|\sigma\right]\right)$ (again, see Appendix A for the definition of $H\left[\xi|\sigma\right]$). The $L^{\text{th}}$ power of $\mathcal{Z}$ and its average then take the form
$\mathcal{Z}^{L}$ $=\sum_{\xi}\prod_{b}\exp\left(\beta\frac{p!}{N^{p-1}}\sum_{a}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right)\,,$
@@
$\approx\frac{1}{2^{MN}}\frac{1}{\left\langle\mathcal{Z}^{*}\right\rangle}\sum_{\xi^{*}}\sum_{\sigma}\exp\Bigg{(}\beta^{*}N\sum_{a\in\Gamma_{*}}\left[\frac{1}{N}\sum_{i}\xi^{*}_{i}\sigma^{a}_{i}\right]^{p^{*}}$
@@ Appendix D Teacher-student free entropy
$\frac{\log\left(Z^{*}\right)}{N}$ =$=\begin{cases}\left(-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}\right)\frac{1}{N}+\log 2\,,&\quad p^{*}=2\,,\\ \frac{1}{2}\left[\beta^{*}\right]^{2}\frac{p^{*}!}{N^{p^{*}-1}}+\log 2+\mathcal{O}\left(\frac{1}{N^{3p^{*}/2-2)}}\right)\,,&\quad p^{*}\geq 3,\end{cases}$
@@
=$=\begin{cases}\frac{1}{2}\left(-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}\right)\alpha^{*}+\log 2\,,&\quad p^{*}=2\,,\\ \frac{1}{2}\left[\beta^{*}\right]^{2}\alpha^{*}+\log 2+\mathcal{O}\left(\frac{1}{N^{p^{*}/2-1}}\right)\,,&\quad p^{*}\geq 3\,,\end{cases}$
@@
$\log\Bigg{\langle}\exp\Bigg{\{}\beta^{*}\frac{p^{*}!}{N^{p^{*}-1}}\sum_{i_{1}<...<i_{p^{*}}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p^{*}}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p^{*}}}+\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\Bigg{\}}\Bigg{\rangle}$
$\quad\approx\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\sum_{b\neq c}\sum_{i_{1}<...<i_{p}}\sum_{j_{1}<...<j_{p}}\xi^{b}_{i_{1}}\xi^{c}_{j_{1}}...\xi^{b}_{i_{p}}\xi^{c}_{j_{p}}\left\langle\sigma^{a}_{i_{1}}\sigma^{a}_{j_{1}}...\sigma^{a}_{i_{p}}\sigma^{a}_{j_{p}}\right\rangle$
$\qquad+\beta^{*}\beta\frac{p^{*}!}{N^{p^{*}-1}}\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p^{*}}}\sum_{j_{1}<...<j_{p}}\left\langle\xi^{*}_{i_{1}}\sigma^{a}_{i_{1}}...\xi^{*}_{i_{p^{*}}}\sigma^{a}_{i_{p^{*}}}\xi^{b}_{j_{1}}\sigma^{a}_{j_{1}}...\xi^{b}_{j_{p}}\sigma^{a}_{j_{p}}\right\rangle$
$\qquad+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN+\frac{1}{2}\left[\beta^{*}\right]^{2}\frac{p^{*}!}{N^{p^{*}-1}}N\,.$
@@
$\log\Bigg{\langle}\exp\Bigg{\{}\beta^{*}\frac{p^{*}!}{N^{p^{*}-1}}\sum_{i_{1}<...<i_{p^{*}}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p^{*}}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p^{*}}}+\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\Bigg{\}}\Bigg{\rangle}$
$\quad=\beta^{2}\frac{p!}{N^{p-1}}N\sum_{b<c}\left[\frac{1}{N}\sum_{i}\xi^{b}_{i}\xi^{c}_{i}\right]^{p}+\beta^{*}\beta\frac{p!}{N^{p-1}}N\sum_{b}\left[\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi^{b}_{i}\right]^{p}$
$\qquad+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN+\frac{1}{2}\left[\beta^{*}\right]^{2}\frac{p!}{N^{p-1}}N\,,$
@@
$\quad\approx\log\Bigg{[}\left\langle\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
$\qquad\left\langle\exp\left\{\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\sigma^{a}_{i_{1}}...\sigma^{a}_{i_{p}}\right\}\right\rangle$
$\qquad+\left\langle\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{j_{1}<...<j_{p}}\xi^{b}_{j_{1}}...\xi^{b}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{j_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle\Bigg{]}\,,$
@@
$\log\left\langle\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle\approx-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}\,,$
@@
$\quad\approx\beta^{2}\frac{p!}{N^{p-1}}N\sum_{b<c}\left[\frac{1}{N}\sum_{i}\xi^{b}_{i}\xi^{c}_{i}\right]^{p}+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN\,,$
@@
$\quad=\left\langle\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{j_{1}<...<j_{p}}\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\xi^{b}_{j_{1}}...\xi^{b}_{j_{p}}\ \xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{j_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
$\quad=\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{j_{1}<...<j_{p}}\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\xi^{b}_{j_{1}}...\xi^{b}_{j_{p}}\left\langle\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{j_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle\,,$
@@
$\quad=\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}$
$\qquad\left\langle\frac{p!}{N^{p}}\sum_{j_{1}<...<j_{p}}\xi^{*}_{j_{1}}...\xi^{*}_{j_{p}}\sigma^{a}_{j_{1}}...\sigma^{a}_{i_{p}}\exp\left\{\beta^{*}\frac{2}{N}\sum_{i_{1}<i_{2}}\xi^{*}_{i_{1}}\xi^{*}_{i_{2}}\sigma^{a}_{i_{1}}\sigma^{a}_{i_{2}}\right\}\right\rangle$
$\quad=V\left(\beta^{*},p\right)\beta\frac{p!}{N^{p-1}}\sum_{b}\sum_{i_{1}<...<i_{p}}\xi^{*}_{i_{1}}...\xi^{*}_{i_{p}}\xi^{b}_{i_{1}}...\xi^{b}_{i_{p}}\,,$
@@
$\quad\approx-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}+\beta^{2}\frac{p!}{N^{p-1}}N\sum_{b<c}\left[\frac{1}{N}\sum_{i}\xi^{b}_{i}\xi^{c}_{i}\right]^{p}+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN$
$\qquad+\left[1-2\beta^{*}\right]^{1/2}\exp\left(\beta^{*}\right)V\left(\beta^{*},p\right)\beta N\sum_{b}\left[\frac{1}{N}\sum_{i}\xi^{*}_{i}\xi^{b}_{i}\right]^{p}\,.$
@@
$=\frac{\left[\beta^{*}\right]^{p/2}}{(p/2)!}\frac{p!}{N^{p/2}}\,,$
@@
$\lambda$ $=\frac{\left[\beta^{*}\right]^{p/2}}{(p/2)!}N^{p/2-1}\sim\mathcal{O}\left(1\right)\,,$
@@
$\quad=\frac{1}{N^{4}}\sum_{j_{1}\neq j_{2}\neq j_{3}\neq j_{4}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\xi^{*}_{j_{3}}\xi^{*}_{j_{4}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\sigma^{a}_{j_{3}}\sigma^{a}_{j_{4}}$
$\quad=\frac{1}{N^{4}}\left[\sum_{j_{1}\neq j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]\left[\sum_{j_{3}\neq j_{4}}\xi^{*}_{j_{3}}\xi^{*}_{j_{4}}\sigma^{a}_{j_{3}}\sigma^{a}_{j_{4}}\right]-\frac{4}{N^{3}}\left[\sum_{j_{1}\neq j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]-\frac{2}{N^{2}}$
$\quad=\frac{1}{N^{2}}\left[\frac{2}{N}\sum_{j_{1}<j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]^{2}-\frac{4}{N^{2}}\left[\frac{2}{N}\sum_{j_{1}<j_{2}}\xi^{*}_{j_{1}}\xi^{*}_{j_{2}}\sigma^{a}_{j_{1}}\sigma^{a}_{j_{2}}\right]-\frac{2}{N^{2}}\,,$
@@
$\frac{\partial\log\left(z^{*}\right)}{\partial\beta^{*}}$ $=\frac{\partial}{\partial\beta^{*}}\left[-\frac{1}{2}\log\left(1-2\beta^{*}\right)-\beta^{*}\right]=\frac{2\beta^{*}}{1-2\beta^{*}}\,,$
@@
$\quad=\prod_{b}\sum_{\xi^{b}_{i}}\exp\left\{\beta k\xi^{b}_{i}\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}\xi^{b}_{i}+x\beta\sqrt{\alpha r}\xi^{b}_{i}\right\}$
$\quad=\prod_{b}\left[2\cosh\left(\beta k\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}+x\beta\sqrt{\alpha r}\right)\right]$
$\quad=2^{L}\cosh^{L}\left(\beta k\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}+x\beta\sqrt{\alpha r}\right)\,,$
@@
$\qquad\exp\Bigg{\{}\beta k\sum_{b}\xi^{b}_{i}\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\sum_{b}\xi^{*}_{i}\xi^{b}_{i}+x\beta\sqrt{\alpha r}\sum_{b}\xi^{b}_{i}\Bigg{\}}\Bigg{\rangle}\Bigg{]}\Bigg{)}\Bigg{\}}$
@@
$\qquad\cosh^{L}\left(\beta k\sigma^{a}_{i}+\beta^{*}\beta\alpha r^{*}\xi^{*}_{i}+x\beta\sqrt{\alpha r}\right)\Bigg{\rangle}\Bigg{]}\Bigg{)}\Bigg{\}}\,.$
@@
$f$ $=\operatorname*{Extr}_{m,k,q,r,q^{*},r^{*}}\Bigg{\{}\beta^{*}\beta\alpha\left[q^{*}\right]^{p}-\frac{1}{2}\beta^{2}\alpha q^{p}+\beta m^{p}-\beta^{*}\beta\alpha r^{*}q^{*}$
@@
$f=\operatorname*{Extr}_{m,k,q^{*},r^{*}}\Bigg{\{}\beta\eta\alpha\left[q^{*}\right]^{p}-\beta m^{p}-\beta\eta\alpha r^{*}q^{*}-\beta mk+\log 2+\Big{\langle}\log\left[\cosh\left(\beta\left[\eta\alpha r^{*}+kz\right]\right)\right]\Big{\rangle}_{z}\Bigg{\}}\,.$